OpenStudy (anonymous):

The function g(t) = 2t represents the number of guitar lessons, g(t), you can complete in t months. How many guitar lessons can you complete in 7 months? Provide your conclusion as a complete sentence.

OpenStudy (campbell_st):

the function is \[g(t) = 2 \times t\] and t is in months so replace t with 7 and calculate the answer

OpenStudy (jordanloveangel):

t is months, and there's 7 months giving so plug that to t in the function of g(t).

OpenStudy (jordanloveangel):

what is the function of g(t) ? it's g(t)=2t plug 7 to t g(7)=2(7)

OpenStudy (anonymous):

So is g(7)=2(7) The answer? @jordanloveangel

OpenStudy (jordanloveangel):

yup ig

OpenStudy (anonymous):

Okay thanks!

OpenStudy (campbell_st):

not quite the answer, you still need to calculate the value

OpenStudy (anonymous):

Oh okay can you explain to me how to do that i am not really good at algebra @campbell_st

OpenStudy (anonymous):

So that would be 14?

OpenStudy (campbell_st):

that's correct

OpenStudy (campbell_st):

so the answer means you would take 14 lessons in the 7 months the function g(t) = 2t is saying you take 2 lessons every month